OpenStudy (petlovers123):

cube with a side length of 4, how can your write the calculations in exponential form?

OpenStudy (petlovers123):

@Mahoganie.Carson

OpenStudy (sshayer):

volume of cube=4^3 surface area of square=6 *4^2

OpenStudy (sshayer):

volume=edge^3 surface area of cube=6*edge^2